# Rectangular recurrence relations in $\mathfrak{gl}_{n}$ and $\mathfrak{o}_{2n+1}$ invariant integrable models

### v1

@@ Abstract.
A new method is introduced to derive general recurrence relations for off-shell Bethe vectors in quantum integrable models with either type $\mathfrak{gl}_{n}$ or type $\mathfrak{o}_{2n+1}$ symmetries. These recurrence relations describe how to add a single parameter, $z$, to specific subsets of Bethe parameters, expressing the resulting Bethe vector as a linear combination of monodromy matrix entries that act on Bethe vectors which do not depend on $z$. We refer to these recurrence relations as rectangular because the monodromy matrix entries involved are drawn from the upper-right rectangular part of the matrix. This construction is achieved within the framework of the zero mode method.
@@ 1 Introduction
One of the fundamental problems in quantum integrable models is the presentation of the Bethe vectors scalar products in a determinant form. For periodic boundary conditions, this problem was investigated and fully solved in [25] for the integrable system associated with the simplest Lie algebra $\mathfrak{gl}_{2}$, where the structure of the Bethe vectors is quite simple.
In the models associated to higher rank symmetries the structure of the Bethe vectors is rather complicated. In principle, the nested Bethe ansatz ensures that the Bethe vectors can be expressed as a combinatorial expression of the monodromy matrix entries acting on a vacuum vector. However, no reasonable formulas for the scalar products can be obtained from this combinatorial presentation.
@@ 2.1 Algebraic context
For any matrix $M\in{\rm End}({\mathbb{C}}^{2n+1})$ we denote by $M^{\rm t}$ the transposition with respect to the secondary diagonal
@@
where the parameters $\chi_{i}$, $i\in I_{\mathfrak{g}}$ are twisting parameters. For $\mathfrak{g}=\mathfrak{gl}_{N}$ these parameters are all independent, while for $\mathfrak{g}=\mathfrak{o}_{2n+1}$ they satisfy the relation $\chi_{i}\chi_{-i}=1$ and $\chi_{0}=1$. For both algebras, the possibility to start the series expansion (2.6) with $\chi_{i}\,\delta_{ij}$ instead of $\delta_{ij}$ corresponds to the usual twist $T(z)\to D\,T(z)$ where $D$ is the diagonal matrix with $D_{ii}=\chi_{i}$. These twisting maps are isomorphisms of the Yangians, since the $R$-matrix commutes with $D\otimes D$ (with the restriction $D^{\rm t}=D$ in the $\mathfrak{o}_{2n+1}$ case).
@@
We describe different embeddings in the Yangians $Y(\mathfrak{gl}_{n})$ and $Y(\mathfrak{o}_{2n+1})$ that will be reflected in a factorisation property of the Bethe vectors, see sections 4.3 and 4.4 below.
@@
Embeddings of $Y(\mathfrak{gl}_{n})\hookrightarrow Y(\mathfrak{o}_{2n+1})$. There are several ways to embed $Y(\mathfrak{gl}_{n})$ in $Y(\mathfrak{o}_{2n+1})$. We describe here four of them that have some relevance for the study of Bethe vectors. Let us denote the $Y(\mathfrak{gl}_{n})$ monodromy matrix embedded in $Y(\mathfrak{o}_{2n+1})$ as ${\mathbb{T}}(z)$.
@@
    $T_{i,j}(z)\equiv{\mathbb{T}}_{i,j}(z)\,,\quad 1\leq i,j\leq n$ (2.13)
@@
  • Another embedding can be done, considering the elements $T_{i,j}(z)$, $-n\leq i,j\leq-1$ and defining ${\mathbb{T}}_{i,j}(z)=T_{-i,-j}(z)$, $1\leq i,j\leq n$. It is easy to see that ${\mathbb{T}}_{i,j}(z)$ also obey the commutation relations (2.5) without the second line. The twisting parameters for the monodromy ${\mathbb{T}}(z)$ in this case are $\chi_{i}^{-1}$.
  • A third embedding can be constructed through the relation (2.9). Indeed, from the relation ${T}(z)=\big{(}T(z+c\kappa_{n})^{-1}\big{)}^{\rm t}$, one deduces that the generators ${\mathbb{T}}_{i,j}(z)=\big{(}T(z+c\kappa_{n})^{-1}\big{)}_{-j,-i}$, for $1\leq i,j\leq n$, also obey the $Y(\mathfrak{gl}_{n})$ commutation relations.
Finally, a fourth embedding can be constructed, that will be used below.
Lemma 2.1.
The monodromy matrix ${\mathbb{T}}(z)$ with matrix entries given by
${\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(z)=-T_{i,j}(z)\,(-1)^{\delta_{i,j}+{{\sf T}_{i,i}/\chi_{i}+{\sf T}_{j,j}/\chi_{j}}},\ \text{with }\begin{cases}{i^{\prime}}=i+n+1,\\ {j^{\prime}}=j+n+1,\end{cases}-n\leq i,j\leq-1\,,$ (2.14)
satisfies the $\mathfrak{gl}_{n}$-type $RTT$ commutation relation with $R$-matrix
$R_{\mathfrak{gl}_{n}}(u,v)=\mathbf{I}\otimes\mathbf{I}+\frac{c\,\mathbf{P}}{u-v}\,.$ (2.15)
Proof.
The monodromy entries $T_{i,j}(u)$ for $-n\leq i,j\leq-1$ of the original $\mathfrak{o}_{2n+1}$ monodromy $T(u)$ satisfy the $\mathfrak{gl}_{n}$-type commutation relations which means that
$[T_{i,j}(u),T_{k,\ell}(v)]=g(u,v)\Big{(}T_{k,j}(v)\ T_{i,\ell}(u)-T_{k,j}(u)\ T_{i,\ell}(v)\Big{)}$
and
$[{{\sf T}_{i,i}},T_{k,\ell}(v)]=\chi_{i}\,(\delta_{i,l}-\delta_{i,k})\ T_{k,\ell}(v)\,.$
The latter equation can be written in the form
$(-1)^{{\sf T}_{i,i}/\chi_{i}}\ T_{k,\ell}(v)\ (-1)^{-{\sf T}_{i,i}/\chi_{i}}=(-1)^{\delta_{i,\ell}}\ (-1)^{\delta_{i,k}}\ T_{k,\ell}(v)\,.$
Using this equality one can verify that the monodromy entries ${\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(u)$ satisfy the $\mathfrak{gl}_{n}$-type commutation relation
$[{\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(u),{\mathbb{T}}_{{k^{\prime}},{\ell^{\prime}}}(v)]=g(u,v)\Big{(}{\mathbb{T}}_{{k^{\prime}},{j^{\prime}}}(v)\ {\mathbb{T}}_{{i^{\prime}},{\ell^{\prime}}}(u)-{\mathbb{T}}_{{k^{\prime}},{j^{\prime}}}(u)\ {\mathbb{T}}_{{i^{\prime}},{\ell^{\prime}}}(v)\Big{)}\,.$
$\square$
Embedding $Y(\mathfrak{o}_{2a+1})\otimes Y(\mathfrak{gl}_{n-a})\hookrightarrow Y(\mathfrak{o}_{2n+1})$. This type of embedding is more intricate. It uses the concept of quasi-determinants [22] $\widehat{T}_{i,j}(z)$ of the monodromy matrix, and we refer to e.g. [23] for more details about quasi-determinants. Indeed, one can show that $\widehat{T}_{i,j}(z)$, $1\leq i,j\leq a$ generate a Yangian subalgebra $Y(\mathfrak{o}_{2a+1})$, while $T_{k+a,\ell+a}(z)$, $1\leq k,\ell\leq n-a$ generate a Yangian subalgebra $Y(\mathfrak{gl}_{n-a})$. Moreover, we have $[\widehat{T}_{i,j}(z)\,,\,T_{k+a,\ell+a}(z)]=0$ for $1\leq i,j\leq a$ and $1\leq k,\ell\leq n-a$. A detailed presentation of this approach is beyond the scope of the present article, we refer to theorem 3.7 and corollary 3.10 in [11] for a detailed presentation of the construction. It is remarkable that this type of embedding can be seen in a simple way on the structure of Bethe vectors, see section 4.4 below.
@@ 2.2 Bethe vectors
with respect of the formal spectral parameter $u$.
Let ${\sf h}_{i}$, $i=1,\ldots,n$ be the operators
@@
for $\mathfrak{o}_{2n+1}$-invariant monodromies.
of Bethe parameters ${\bar{t}}^{s}=\{{\bar{t}}^{s}_{1},\ldots,{\bar{t}}^{s}_{r_{s}}\}$ with cardinality $|{\bar{t}}^{s}|=r_{s}$. The superscript on the (sets of) Bethe parameters denotes the color of the (set of) Bethe parameters. The colors are in correspondence with the simple roots of the algebra $\mathfrak{g}$. The Bethe vectors themselves are certain polynomials of the monodromy entries $T_{i,j}(u)$ for $i\leq j$ acting to the right vacuum vector $|0\rangle$
${\mathbb{B}}(\bar{t})=\mathcal{P}\big{(}T_{i\leq j}(\bar{t})\big{)}\ |0\rangle=\mathcal{B}(\bar{t})\ |0\rangle,$ (2.26)
where the polynomial $\mathcal{B}(\bar{t})=\mathcal{P}\big{(}T_{i\leq j}(\bar{t})\big{)}$ is called a pre-Bethe vector.
Analogously, left or dual Bethe vectors are polynomials of monodromy entries $T_{i,j}(u)$ for $i\geq j$ acting to the left vacuum vector $|0\rangle$
${\mathbb{C}}(\bar{t})=\langle 0|\ \mathcal{P}(T_{i\geq j}(\bar{t}))=\langle 0|\ \mathcal{C}(\bar{t}).$ (2.27)
@@
The off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$ and ${\mathbb{C}}({\bar{t}})$ are eigenvectors of the operators ${\sf h}_{i}$ with eigenvalues
${\sf h}_{i}\cdot{\mathbb{B}}({\bar{t}})=(r_{i-1}-r_{i})\ {\mathbb{B}}({\bar{t}}),\qquad{\mathbb{C}}({\bar{t}})\cdot{\sf h}_{i}=(r_{i-1}-r_{i})\ {\mathbb{C}}({\bar{t}}),\qquad i=1,\ldots,n.$ (2.28)
@@
if each set ${\bar{t}}^{s}$ of Bethe parameters satisfy the equalities
$\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})=\frac{\lambda_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{\lambda_{s+1}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})}=\frac{f_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})}{f_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}\ \frac{f({\bar{t}}^{s+1},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{f({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1})}\quad\text{with}\begin{cases}s=1,\ldots,n-1,\quad\text{for }\ \mathfrak{gl}_{n},\\ s=0,1,\ldots,n-1,\quad\text{for }\ \mathfrak{o}_{2n+1},\end{cases}$ (2.34)
@@ 3.1 Action formula for $T_{1,n}(z)$ and ${\sf T}_{\ell+1,\ell}$
$\begin{split}{\sf T}_{\ell+1,\ell}\cdot{\mathbb{B}}({\bar{t}})&=\sum_{\rm part}\Big{(}\chi_{\ell+1}\ \alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})-\\ &\quad-\chi_{\ell}\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\Big{)}\ {\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\,,\end{split}$ (3.7)
@@ 3.2 Rectangular recurrence relations for $\mathfrak{gl}_{n}$
$\Xi^{\ell,k}_{i,j}(z;{\bar{t}})=\psi_{\ell}(z;{\bar{t}}_{\scriptscriptstyle\rm I})\,\phi_{k}(z;{\bar{t}}_{\scriptscriptstyle\rm III})\prod_{s=i}^{\ell-1}\Omega^{R}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s=k}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\ \Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})$ (3.11)
@@
The sets ${\bar{t}}^{\ell},\ldots,{\bar{t}}^{k-1}$ are not parted and in (3.10) ${\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{\ell}$ and ${\bar{t}}^{k-1}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{k-1}$. If the set ${\bar{t}}^{s}=\varnothing$ and according to (3.12) one may have $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=1$ for some $i$ or $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=1$ for some $j$, then the terms in the r.h.s. of (3.10) for such values of the indices $i$ and $j$ have to be discarded.
@@
The proposition 3.1 will be proved in appendix A. $\square$
@@
$\begin{split}{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{{\bar{t}}^{\ell},z\},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})&=\sum_{i=1}^{\ell}\sum_{j=\ell+1}^{n}\sum_{\rm part}\frac{T^{+}_{i,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},{\bar{t}}^{\ell},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{\ell+1}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})}{\lambda_{\ell+1}(z)\,g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II})h(z,{\bar{t}}^{\ell})h({\bar{t}}^{\ell},z)g({\bar{t}}^{\ell+1}_{\scriptscriptstyle\rm II},z)}\times\\ &\quad\times\prod_{s=i}^{\ell-1}\gamma({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})\frac{h({\bar{t}}^{s+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II})}\prod_{s=\ell+1}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\gamma({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm III},\bar{t}^{s-1}_{\scriptscriptstyle\rm II})}{g({\bar{t}}^{s+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})}\,.\end{split}$

### v2

# LAPTH-070/24

@@ Abstract.
A new method is introduced to derive general recurrence relations for off-shell Bethe vectors in quantum integrable models with either type $\mathfrak{gl}_{n}$ or type $\mathfrak{o}_{2n+1}$ symmetries. These recurrence relations describe how to add a single parameter $z$, to specific subsets of Bethe parameters, expressing the resulting Bethe vector as a linear combination of monodromy matrix entries that act on Bethe vectors which do not depend on $z$. We refer to these recurrence relations as rectangular because the monodromy matrix entries involved are drawn from the upper-right rectangular part of the matrix. This construction is achieved within the framework of the zero mode method.
@@ 1 Introduction
One of the key problems in quantum integrable models is the presentation of the Bethe vectors scalar products in a determinant form. For periodic boundary conditions, this problem was investigated and fully solved in [26] for the integrable system associated with the simplest Lie algebra $\mathfrak{gl}_{2}$, where the structure of the Bethe vectors is quite simple.
In the models associated to higher rank symmetries the structure of the Bethe vectors is rather complicated. In principle, the nested Bethe ansatz ([16, 17] for $\mathfrak{gl}_{n}$ and [25] for $\mathfrak{o}_{n}$) ensures that the Bethe vectors can be expressed as a combinatorial expression of the monodromy matrix entries acting on a vacuum vector. However, no reasonable formulas for the scalar products can be obtained from this combinatorial presentation.
@@ 2.1 Algebraic context
In the $\mathfrak{o}_{2n+1}$-case for any matrix $M\in{\rm End}({\mathbb{C}}^{2n+1})$ we denote by $M^{\rm t}$ the transposition with respect to the secondary diagonal
@@
where the parameters $\chi_{i}$, $i\in I_{\mathfrak{g}}$ are twisting parameters. For $\mathfrak{g}=\mathfrak{gl}_{N}$ these parameters are all independent, while for $\mathfrak{g}=\mathfrak{o}_{2n+1}$ they satisfy the relation $\chi_{i}\chi_{-i}=1$ and $\chi_{0}=1$. For both algebras, the possibility to start the series expansion (2.6) with $\chi_{i}\,\delta_{ij}$ instead of $\delta_{ij}$ corresponds to the usual twist $T(z)\to D\,T(z)$ where $D$ is the diagonal matrix with $D_{ii}=\chi_{i}$. These twisting maps are isomorphisms between two RTT-algebras, since the $R$-matrix commutes with $D\otimes D$ (with the restriction $D^{\rm t}=D$ in the $\mathfrak{o}_{2n+1}$ case).
@@
We describe different embeddings in the Yangians $Y(\mathfrak{gl}_{n})$ and $Y(\mathfrak{o}_{2n+1})$ that will be reflected in a factorisation (also called splitting) property of the Bethe vectors, see sections 4.3 and 4.4 below.
@@
Embeddings of $Y(\mathfrak{gl}_{n})\hookrightarrow Y(\mathfrak{o}_{2n+1})$. There are several ways to embed $Y(\mathfrak{gl}_{n})$ in $Y(\mathfrak{o}_{2n+1})$. We describe here two of them that have some relevance for the study of Bethe vectors. Let us denote the $Y(\mathfrak{gl}_{n})$ monodromy matrix embedded in $Y(\mathfrak{o}_{2n+1})$ as ${\mathbb{T}}(z)$.
@@
    ${\mathbb{T}}_{i,j}(z)\equiv T_{i,j}(z)\,,\quad 1\leq i,j\leq n$ (2.13)
@@
  • Another embedding can be done, considering the elements $T_{i,j}(z)$, $-n\leq i,j\leq-1$ and defining
    $\hat{{\mathbb{T}}}_{i,j}(z)=T_{i-n-1,j-n-1}(z),\qquad 1\leq i,j\leq n.$ (2.14)
    It is easy to see that $\hat{{\mathbb{T}}}_{i,j}(z)$ also obey the commutation relations (2.5) without the second line. The twisting parameters for the monodromy $\hat{{\mathbb{T}}}(z)$ in this case are $(\chi_{n+1-i})^{-1}$.
Embedding $Y(\mathfrak{o}_{2a+1})\otimes Y(\mathfrak{gl}_{n-a})\hookrightarrow Y(\mathfrak{o}_{2n+1})$. This type of embedding is more intricate. It uses the concept of quasi-determinants [4] $\widehat{T}_{i,j}(z)$ of the monodromy matrix, and we refer to e.g. [23] for more details about quasi-determinants in the Yangian case. Indeed, one can show that $\widehat{T}_{i,j}(z)$, $1\leq i,j\leq a$ generate a Yangian subalgebra $Y(\mathfrak{o}_{2a+1})$, while $T_{k+a,l+a}(z)$, $1\leq k,l\leq n-a$ generate a Yangian subalgebra $Y(\mathfrak{gl}_{n-a})$. Moreover, we have $[\widehat{T}_{i,j}(z)\,,\,T_{k+a,l+a}(z)]=0$ for $1\leq i,j\leq a$ and $1\leq k,l\leq n-a$. A detailed presentation of this approach is beyond the scope of the present article, we refer to theorem 3.7 and corollary 3.10 in [12] for a detailed presentation of the construction. It is remarkable that this type of embedding is reflected in a simple way on the structure of Bethe vectors, see section 4.4 below.
@@ 2.2 Bethe vectors
with respect of the formal parameter $z$.
of Bethe parameters ${\bar{t}}^{s}=\{{\bar{t}}^{s}_{1},\ldots,{\bar{t}}^{s}_{r_{s}}\}$ with cardinality $|{\bar{t}}^{s}|=r_{s}$. The superscript on the (sets of) Bethe parameters denotes its color. The colors are in correspondence with the simple roots of the algebra $\mathfrak{g}$. This can be formalized through the operators
@@
for $\mathfrak{o}_{2n+1}$-invariant monodromies. Using the operators ${\sf h}_{i}$ (2.21) one also defines the operators
$\mathsf{t}_{s}=\sum_{i=s+1}^{n}\ {\sf h}_{i}\,,$ (2.25)
where $s=1,\ldots,n-1$ for $\mathfrak{gl}_{N}$-invariant monodromies and $s=0,1,\ldots,n-1$ for $\mathfrak{o}_{2n+1}$-invariant monodromies.
The Bethe vectors themselves are certain polynomials of the monodromy entries $T_{i,j}(u)$ for $i<j$ depending on various Bethe parameters acting to the right vacuum vector $|0\rangle$
${\mathbb{B}}(\bar{t})=\mathcal{P}\big{(}T_{i<j}(\bar{t})\big{)}\ |0\rangle=\mathcal{B}(\bar{t})\ |0\rangle,$ (2.26)
where the polynomial $\mathcal{B}(\bar{t})=\mathcal{P}\big{(}T_{i<j}(\bar{t})\big{)}$ is called a pre-Bethe vector.
Analogously, left or dual Bethe vectors are polynomials of monodromy entries $T_{i,j}(u)$ for $i>j$ acting to the left vacuum vector $|0\rangle$
${\mathbb{C}}(\bar{t})=\langle 0|\ \mathcal{P}(T_{i>j}(\bar{t}))=\langle 0|\ \mathcal{C}(\bar{t}).$ (2.27)
@@
The off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$ and ${\mathbb{C}}({\bar{t}})$ are eigenvectors of the operators $\mathsf{t}_{s}$ (2.25):
$\mathsf{t}_{s}\cdot{\mathbb{B}}({\bar{t}})=r_{s}\ {\mathbb{B}}({\bar{t}}),\qquad{\mathbb{C}}({\bar{t}})\cdot\mathsf{t}_{s}=r_{s}\ {\mathbb{C}}({\bar{t}}),\qquad i=1,\ldots,n.$ (2.28)
with $s=1,\ldots,n-1$ for $\mathfrak{gl}_{n}$ and $s=0,1,\ldots,n-1$ for $\mathfrak{o}_{2n+1}$. This property can be proved using the recurrence relations for the Bethe vectors and the action of the monodromy matrix entries $T_{i,i}(z)$ on Bethe vectors.
@@
if each set ${\bar{t}}^{s}$ of Bethe parameters satisfy the Bethe equations
$\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})=\frac{\lambda_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{\lambda_{s+1}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})}=\frac{f_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})}{f_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}\ \frac{f({\bar{t}}^{s+1},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{f({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1})}\quad\text{with }\begin{cases}s=1,\ldots,n-1,\quad&\text{for }\ \mathfrak{gl}_{n},\\ s=0,1,\ldots,n-1,\quad&\text{for }\ \mathfrak{o}_{2n+1},\end{cases}$ (2.34)
@@ 3.1 Action formula for $T_{1,n}(z)$ and ${\sf T}_{\ell+1,\ell}$
$\begin{split}{\sf T}_{\ell+1,\ell}\cdot{\mathbb{B}}({\bar{t}})=\sum_{\rm part}&\Big{(}\chi_{\ell+1}\ \alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})-\\ &\quad-\chi_{\ell}\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\Big{)}\ {\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\,,\end{split}$ (3.7)
@@ 3.2 Rectangular recurrence relations for $\mathfrak{gl}_{n}$
$\Xi^{\ell,k}_{i,j}(z;{\bar{t}})=g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{k}_{\scriptscriptstyle\rm III},z)\prod_{s=i}^{\ell-1}\Omega^{R}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s=k}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\ \Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})$ (3.11)
@@
The sets ${\bar{t}}^{\ell},\ldots,{\bar{t}}^{k-1}$ are not parted and in (3.10) ${\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{\ell}$ and ${\bar{t}}^{k-1}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{k-1}$.
If the set ${\bar{t}}^{s}=\varnothing$ and according to (3.12) one may have $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=1$ for some $i$ or $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=1$ for some $j$, then the terms in the r.h.s. of (3.10) for such values of the indices $i$ and $j$ have to be discarded.
@@
The proposition 3.1 will be proved in appendix A by induction starting with (3.5). $\square$
@@
$\begin{split}{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{{\bar{t}}^{\ell},z\},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})&=\sum_{i=1}^{\ell}\sum_{j=\ell+1}^{n}\sum_{\rm part}\frac{T_{i,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},{\bar{t}}^{\ell},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{\ell+1}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})}{\lambda_{\ell+1}(z)\,g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II})h(z,{\bar{t}}^{\ell})h({\bar{t}}^{\ell},z)g({\bar{t}}^{\ell+1}_{\scriptscriptstyle\rm II},z)}\times\\ &\quad\times\prod_{s=i}^{\ell-1}\gamma({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})\frac{h({\bar{t}}^{s+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II})}\prod_{s=\ell+1}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\gamma({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm III},\bar{t}^{s-1}_{\scriptscriptstyle\rm II})}{g({\bar{t}}^{s+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})}\,.\end{split}$
Here the sum over partition is the same as in the proposition 3.1.